# TopoGEN: topology-driven microstructure generation for in silico modeling of fiber network mechanics

### v1

@@ 2.3 Case study: rheological tests on collagen networks
We assess the role of microstructural features, including average valency, average fiber length and thickness, fiber elastic modulus, and network concentration on the differential elastic modulus, defined as the derivative of the stress with respect to the strain, as in Jansen et al. (2018). Our analysis employs the first Piola–Kirchhoff stress measure $\mathbf{P}$ (Equation 2), which is first smoothed and then differentiated with respect to nominal strain to yield a nominal differential shear modulus. We compare our results with experimental data taken from Jansen et al. (2018), where the authors investigated (i) the nonlinear elastic behavior of rat tail collagen type I networks polymerized at different temperatures using a rheometer and (ii) network architecture features, including valency, fibril diameter and length through multiple imaging modalities. Briefly, Jansen and coauthors found that higher polymerization temperatures reduce the average valency and the fibril diameter (Table 6.1). We virtually replicate these polymerized networks and their respective microstructural architectures by matching the in silico collagen concentration ($\rho_{c}$) and the experimental values, as in Eichinger et al. (2021a), through:
@@
where $R_{f}$ is the fiber radius, $L_{tot}$ represents the sum of all individual fiber lengths, $V_{RVE}$ denotes the volume of the RVE, and $v_{c}$ is the specific volume of collagen fibers (Hulmes and Miller, 1979).
@@ 3.1 Generation of microstructure-informed fiber networks
Figure 3 highlights the power of our TopoGEN pipeline to generate various topology-informed discrete fiber networks. Without loss of generality, we here leverage TopoGEN to generate collagen type I hydrogels representing the experimental in vitro networks generated by Jansen et al. (2018). Starting from random network initializations, we use simulated annealing to align the network architecture in terms of valency and edge length distributions, prioritized as key parameters for capturing isotropic network properties. To match the valency observed in self-assembled collagen type-I networks, where fibers form branches (valency 3) or cross-links (valency 4), we constrain our network to accept only these valency values. In particular, our generated discrete fiber network topologies replicate the observed decrease in average network valency associated with increasing temperatures. Therefore, as the temperature rises, inter-fiber interactions shift towards a prevalence of branch points over cross-links as can be seen from the decreasing valency in the top-left subplot of Figure 3. The length optimization follows a two-step process. First, we remove small, unrepresentative fiber segments during the valency optimization stage to ensure a physically meaningful network structure. To achieve this, we assign higher removal probabilities to shorter edges and iteratively optimize valency following this weighted edge removal approach. Next, we adjust node positions to align with the target length log-normal distribution using simulated annealing guided by the KL divergence metric. This metric demonstrates superior performance, achieving more significant improvement within runtimes comparable to a conventional technique, such as the Cramér-von Mises (CVM) test (Lindström et al. (2010, 2013); Eichinger et al. (2021a)). See Table S.2 for a detailed comparison across 20 independent tests of the performance metrics between the two methods. The average fiber length is optimized to match the experimentally measured fiber lengths observed across different temperatures (Table 6.1), and the resulting log-normal distribution is depicted in the top-right subplot of Figure 3. In the bottom row of Figure 3, two-dimensional projections of our three-dimensional in silico reconstructions are compared to high-resolution images of reconstituted collagen I network, taken from Jansen et al. (2018).
@@
Figure 3: TopoGEN in action, experimentally informed discrete fiber network topology optimization. The top row shows simulated annealing results of the valency and length optimization process for different temperatures. Increasing the temperature decreases the network’s average valency ($\overline{z}$), leading to more branches forming between fibers (left plot). Weighted edge removal facilitates an intermediate optimization stage, where shorter edges are preferentially removed during system dilution. Subsequently, the algorithm shifts the intermediate state toward the final optimized configuration by rearranging the network’s nodes. The length optimization is shown here for the 37 °C condition at a concentration of 4 mg/mL (central plot). The four images on the bottom row compare scanning electron microscopy (SEM) images (taken from Jansen et al. (2018)) of reconstituted collagen I network in vitro (central images) with the simulated maximum intensity projection of our three-dimensional in silico reconstructions (external images) at 30 °C and 37 °C, respectively. The black squares in the experimental and simulation images have an edge length of 20 µm.
@@ 3.2 Boundary value problem at the mesoscale
Figure 4 showcases an exemplary finite element simulation outcome of a representative fiber network generated at 37 °C, with an average valency of 3.12, a concentration of 4 mg/mL, and an average fiber length of 2 µm. The generated RVE undergoes uniaxial, biaxial, and simple shear deformations with fully periodic boundary conditions. We conduct all our RVE mechanical deformation analyses using Abaqus/Standard. To ensure convergence, we adhere to Abaqus’ default convergence criteria and employ automatic stabilization. We strictly require that at every increment, the ratio of viscous damping energy (reported as ALLSD) to total strain energy (reported as ALLSE) does not exceed 2% (Kakaletsis et al., 2023).
Figure 4: Multiaxial loading with periodic boundary conditions. Fiber logarithmic strain distribution in a representative (37 °C, average valency = 3.12, concentration = 4 mg/mL, average fiber length = 2 µm) RVE under uniaxial (left), biaxial (middle), and simple shear (right) loading, with 50% strain applied for each loading condition. The undeformed cubic domain is shown with dashed lines and the deformed domain with solid lines.
@@ 3.3 Microstructure-function relationships in simulated collagen networks
To evaluate how each microstructural feature influences the network’s differential shear modulus, we systematically vary each structural parameter based on experimental observations for collagen type I networks (Table 6.1) and perform simple shear tests up to 50% strain on ten representative samples per tested topological or microstructural alteration. We analyze the impact that these topological and microstructural changes have on the differential shear modulus across increasing strain levels relative to the average value of the original network architecture. Figure 5 illustrates the mechanical response for various topological and microstructural parameters derived from experiments and our simulations. We examine the role of additional network-level factors (i.e., valency and concentration) and fibril-level characteristics (i.e., radius, stiffness, and average length) using our 3D simulations. We systematically alter these parameters within the experimentally reported ranges from Jansen et al. (2018) (Table 6.1) for diameter and average fiber length, and from Mahutga et al. (2023); Gacek et al. (2023) for fibril stiffness. We find that transitioning from the low-strain to the high-strain regime results in a nonlinear stiffening for all tested topologies. In particular, for all collagen type I networks, we observe a two-order increase in strain stiffening from the initial to the maximum differential shear modulus, with significant stiffening emerging in the medium- to high-strain regime. Our model results are consistent with the experimental data from Jansen et al. (2018) where the concentration is fixed at 4 mg/mL while the average valency is varied based on experimental measurements (Figure 5 - experimental data and valency effect). Conversely, variations in concentration exhibit an uneven impact, with the most significant changes in the differential shear modulus occurring in the low-strain regime (Figure 5 - concentration effect). In this regime, fibrils primarily undergo non-affine rotations to align with the principal loading direction. Notably, increasing the concentration from 1.5 to 4 mg/mL results in an order-of-magnitude increase in the initial differential shear modulus. In contrast, within the large-strain regime, the maximum modulus remains largely insensitive to concentration changes. At a concentration of 1.5 mg/mL, our models predict an almost three-order-of-magnitude increase in modulus when transitioning from the low- to high-strain regime for both valencies tested.
@@
Figure 5: Analysis of microstructure-mechanics relationship in reconstituted collagen networks. The differential shear modulus is shown as a function of strain on a log-log scale for each tested condition. The top row presents (1) experimental results from rheological tests on reconstituted collagen networks with a concentration of 4 mg/mL at varying average valencies (Jansen et al., 2018) and our simulation results from shear tests (up to 50% strain) for (2) 4 mg/mL samples at the same valencies tested experimentally and (3) decreasing concentrations at fixed valencies of 3.12 and 3.56. In all the top row simulations, the fibril-related parameters are set to their minimum values, with fibril radius of 80 nm, Young’s modulus of 700 MPa in tension and 70 MPa in compression, and an average length of 2 µm. The bottom row shows simulation results under the same loading conditions as the top row, focusing on the impact of collagen fibril morphology and micromechanics. These results examine inter- and intra-valency effects at a 4 mg/mL fixed concentration. All simulation results represent the response of ten samples per condition, with the average response shown as a line and the standard deviation indicated by the shaded area.
With regard to the morphological changes of the fibers, variations in radius and length do not exhibit noticeable effects across all tested valencies and strain ranges (Figure 5 - fibril radius and length effects). Overall, increasing the radius slightly increases the network’s differential shear modulus, whereas increasing the length leads to its reduction. This is because longer, more slender fibrils are more susceptible to bending, particularly in the low-strain regime when they are not aligned with the primary loading direction. Increasing Young’s modulus from 700 MPa to 1 GPa (Figure 5 - fibril stiffness effects) yields comparable behavior for the medium-to-high-strain regime. However, it exerts a more pronounced influence in the low-strain regime. Notably, these findings remain consistent across different valency conditions, suggesting that they are independent of the polymerization temperature used to form the network. Table 1 summarizes the deformation-dependent influence of each microstructural parameter on the network’s bulk elastic properties. To build the table, we compare the variation in the differential shear modulus for each microstructural parameter by averaging data from 10 independent samples for each tested condition. At predetermined strain points, we compute the mean and standard deviation for each condition and derive the normalized difference between the means. We aggregate these normalized changes across all parameters and compute the 33rd and 67th percentiles to set thresholds: values below the 33rd percentile indicate a low impact, values between the 33rd and 67th percentiles indicate a moderate impact, and values above the 67th percentile indicate a high impact.
@@
strain [-] 0.010 0.125 0.250 0.375 0.500
[…]
fibrils radius $\approx$ $\approx$ $\uparrow$ $\approx$ $\approx$
[…]
fibrils length $\approx$ $\approx$ $\downarrow$ $\approx$ $\approx$
@@ 3.4 Non-affinity in simulated collagen network
Figure 6 illustrates the discrepancy between the nodal displacements predicted by our in silico simulations and the affine predictions based on the deformation gradient imposed at the boundaries. Here, we focus on two specific cases: loosely connected networks with an average valency of $\overline{z}$ = 3.12, and more interconnected networks with average valency of $\overline{z}$ = 3.56. We define the non-affinity score ($NS$) as the normalized root mean squared error between purely affine deformation ($u_{affine}$) of the whole network (i.e., the displacements obtained by interpolating the macroscopic deformation gradient imposed at the boundaries to all internal nodes) and the displacements predicted in our simulations ($u_{predicted}$):
@@
where $n$ is the total number of nodes.
Figure 6: Non-affinity score with nodal displacements. Comparison of predicted affine displacements (gray) and computed displacements for five samples, with average valency $\overline{z}=$ 3.12 (top row) and valency $\overline{z}=$ 3.56 (bottom row), at three distinct strain levels. For low-strain regime non-affine modes predominate, with a large set of nodes not following the boundary displacement based on the macroscopic deformation gradient. At strain 0.5, strain-induced affinity enhances the agreement between predicted and measured distributions.
The networks exhibit different behavior from the affine model under all the tested conditions. The most significant discrepancies occur in the low- to medium-strain regime, where the $NS$ score decreases fivefold. In contrast, the medium- to high-strain regime experiences only a twofold reduction. In the low-strain regime, very few nodes adhere to the boundary displacement prescribed by the imposed macroscopic deformation gradient. However, a noticeable shift occurs in the medium- to high-strain regime, where nodal displacements closely follow an overall affine displacement pattern. Both loosely connected (with $\overline{z}=$ 3.12) and more interconnected (with $\overline{z}=$ 3.56) networks converge toward a more affine response at higher strain levels. Networks with lower valency exhibit slightly increased non-affinity in the low-strain regime than those with higher valency. With fewer fibers aligning with the affine deformation, load distribution in these low-valencies, loosely connected networks, becomes highly localized. Increasing the number of cross-links compared to the branches (with $\overline{z}=$ 3.56) results in a marginally improved match between the theoretical affine predictions and the experimental measurements. Overall, these sub-static networks exhibit nodal deformation patterns that, in the low- to medium-strain regime, deviate from affine deformation assumptions.
@@ 4.1 Topological complexity and heterogeneity
Many materials - not just biological materials - are made of fibers. Natural fibrous materials include cotton, wool, and silk, while industrial examples range from polymeric foams and rubber to felt and paper. Among biological materials, an important fibrous system is the extracellular matrix, where collagen fibers assemble into networks to bear mechanical loads and transmit mechanical information over long distances (Ban et al., 2016). Another crucial fibrous system is fibrin, derived from fibrinogen, which assembles into a fibrillar network upon tissue injury to form blood clots. These clots must withstand high tensile forces from blood flow while ensuring rapid degradation after wound healing to prevent thrombosis (Pieters and Wolberg, 2019). Accurately modeling this seemingly passive fibrous environment through computational methods remains a significant challenge. A key requirement is to account for the three-dimensional arrangement of these networks and their structural heterogeneity, which affects the local cellular environment and influences the overall mechanical response of tissues. In this study, we employ a mechanistic approach that explicitly models the interconnected fiber networks underlying soft tissues by leveraging high-resolution imaging data. This framework enables us to investigate how macroscopic material behavior emerges from its microscopic constituents, accounting for microstructural variations. Our generative pipeline builds on the framework by Lindström et al. (2010), where high-resolution confocal images are adopted to derive an image-based representation of collagen networks using simulated annealing. The iterative optimization process is designed to transform initially random Voronoi networks into structures that accurately represent the target biological network in terms of valency and length distribution. Compared to Lindström et al. (2010), we capture the non-affine localized effects by restricting the network to include only cross-links and branches. For the length optimization, we employ Kullback-Leibler divergence to compare the actual and target distributions throughout the iterative optimization process. As a result, we improve the energy accuracy of the optimization process without increasing runtimes compared to the commonly used (Lindström et al., 2010, 2013; Nan et al., 2018; Eichinger et al., 2021a) CVM tests. Notably, while the probabilistic search strategy of simulated annealing cannot guarantee convergence of the optimization scheme to the global minimum (Suman and Kumar, 2006), it is sufficient for our purposes that the final configuration statistically aligns with experimental measurements without requiring exact correspondence. While focused on collagen-like networks, our modeling pipeline can adapt to other biomaterials and topologies. For instance, we can effectively tune the fibers to replicate the morphology and mechanics of fibrin networks to simulate the mechanical behavior of another essential load-bearing biopolymer in the extracellular space. Alternatively, we can introduce higher valencies by merging nodes within a specific distance threshold, or we can disregard the bending rigidity to model fibers as springs transmitting only central forces, as observed in the intracellular environment of semi-flexible F-actin networks and intermediate filaments (Licup et al., 2015). Leveraging detailed microstructural information, our approach can be extended to model non-living networks. For example, with random Voronoi seeds, we can replicate the microstructure of paper while with regularly arranged seeds, we can create ordered structures, such as those found in fabrics (Picu, 2011).
@@
4.2 Impact of microstructure on stress-strain behavior

### v2

@@ 2.3 Case study: rheological tests on collagen networks
We evaluate the role of microstructural features, including average connectivity, average fiber length and thickness, fiber elastic modulus, and network concentration, on the differential elastic modulus, defined as the derivative of the stress with respect to the shear strain, as in [43]. Our analysis employs the first Piola–Kirchhoff stress measure $\mathbf{P}$ (Equation 2), which is first smoothed and then differentiated with respect to RVE shear to yield a nominal differential shear modulus. We compare our results with experimental data taken from [43], where the authors investigated (i) the nonlinear elastic behavior of rat tail collagen type I networks polymerized at different temperatures using a rheometer and (ii) network architecture features, including connectivity, fibril diameter, and length through multiple imaging modalities. Briefly, Jansen and coauthors found that higher polymerization temperatures reduce the average connectivity and the fibril diameter (Table 6.1). We virtually replicate these polymerized networks and their respective microstructural architectures by matching the in silico collagen concentration ($\rho_{c}$) and the experimental values, as in [24], through:
@@
where $R_{f}$ is the fiber radius, $L_{tot}$ represents the sum of all individual fiber lengths, $V_{RVE}$ denotes the volume of the RVE, and $v_{c}$ is the specific volume of collagen fibers [37].
@@ 3.1 Generation of microstructure-informed fiber networks
Figure 3 highlights the power of our TopoGEN pipeline to generate various topology-informed discrete fiber networks.
@@
Figure 3: TopoGEN in action, experimentally informed discrete fiber network topology optimization. The top row shows two-step connectivity and length optimization results across temperatures. Increasing the temperature decreases the network’s average connectivity ($\overline{z}$), leading to more branches forming between fibers (left plot). Weighted edge removal facilitates an intermediate optimization stage, where shorter edges are preferentially removed during system dilution. Subsequently, the algorithm shifts the intermediate state toward the final optimized configuration by rearranging the network’s nodes. The length optimization is shown here for the 37 °C condition at a concentration of 4 mg/mL (central plot). The four images on the bottom row compare scanning electron microscopy (SEM) images (taken from [43]) of reconstituted collagen I network in vitro (central images) with the simulated maximum intensity projection of our three-dimensional in silico reconstructions (external images) at 30 °C and 37 °C, respectively. The black squares in the experimental and simulation images have an edge length of 20 µm.
Without loss of generality, we here leverage TopoGEN to generate collagen type I hydrogels representing the experimental in vitro networks generated by [43]. Starting from random network initializations, we use simulated annealing to align the network architecture in terms of connectivity and edge length distributions, prioritized as key parameters for capturing isotropic network properties. To match the connectivity observed in self-assembled collagen type-I networks, where fibers form branches (connectivity 3) or cross-links (connectivity 4), we constrain our network to accept only these connectivity values. In particular, our generated discrete fiber network topologies replicate the observed decrease in average cross-link density associated with increasing temperatures. Therefore, as the temperature rises, inter-fiber interactions shift towards a prevalence of branch points over cross-links, as can be seen from the decreasing connectivity in the top-left subplot of Figure 3. The length optimization follows a two-step process. First, we remove small, unrepresentative fiber segments during the connectivity optimization stage to ensure a physically meaningful network structure. To achieve this, we assign higher removal probabilities to shorter edges and iteratively optimize connectivity following this weighted edge removal approach. Next, we adjust node positions to align with the target length log-normal distribution using simulated annealing guided by the KL divergence metric. This metric demonstrates superior performance, achieving more significant improvement within runtimes comparable to a conventional technique, such as the Cramér-von Mises (CVM) test ([53, 52, 24]). See Table S.2 for a detailed comparison across 20 independent tests of the performance metrics between the two methods. The average fiber length is optimized to match the experimentally measured fiber lengths observed across different temperatures (Table 6.1), and the resulting log-normal distribution is depicted in the top-right subplot of Figure 3. In the bottom row of Figure 3, two-dimensional projections of our three-dimensional in silico reconstructions are compared to high-resolution images of reconstituted collagen I network, taken from [43]. For the illustrated topology, the optimization convergence is achieved within $10^{5}$ number of iterations in 81 seconds on a 64-bit system with a 12th Gen Intel(R) Core(TM) i9-12900H @ 2.50 GHz and 32.0 GB RAM. See Figure S.2 for details on the optimization performance.
@@ 3.2 Boundary value problem at the mesoscale
Figure 4 showcases an exemplary finite element simulation outcome of a representative fiber network generated at 37 °C, with an average connectivity of 3.12, a concentration of 4 mg/mL, and an average fiber length of 2 µm.
Figure 4: Multi-axial loading with periodic boundary conditions. Fiber logarithmic stretch distribution in a representative (37 °C, average connectivity 3.12, concentration 4 mg/mL, average fiber length 2 µm) RVE under uniaxial (left), biaxial (middle), and simple shear (right) loading, with 50% stretch applied for each loading condition. The undeformed cubic domain is shown with dashed lines and the deformed domain with solid lines.
The generated RVE undergoes uniaxial, biaxial, and simple shear deformations with fully periodic boundary conditions. We conduct all our RVE mechanical deformation analyses using Abaqus/Standard. The average number of degrees of freedom in our simulations is $5\times 10^{4}$. To ensure convergence, we adhere to Abaqus’ default convergence criteria and employ automatic stabilization. We strictly require that at every increment, the ratio of viscous damping energy (reported as ALLSD) to total strain energy (reported as ALLSE) does not exceed 2% [69, 44].
@@ 3.3 Microstructure-function relationships in simulated collagen networks
To evaluate how each microstructural feature influences the network’s differential shear modulus, we systematically vary each structural parameter based on experimental observations for collagen type I networks (Table 6.1) and perform simple shear tests up to 50% shear on ten representative samples per tested topological or microstructural alteration. We analyze the impact that these topological and microstructural changes have on the differential shear modulus across increasing stretch levels relative to the average value of the original network architecture.
@@
Figure 5: Analysis of microstructure-mechanics relationship in reconstituted collagen networks. The differential shear modulus is shown as a function of rescaled stretch $(\lambda-1)$ on a log-log scale for each tested condition. The top row presents (1) experimental results from rheological tests on reconstituted collagen networks with a concentration of 4 mg/mL at varying average connectivity [43] and our simulation results from shear tests (up to 50% shear) for (2) 4 mg/mL samples at the same connectivity tested experimentally and (3) decreasing concentrations at fixed connectivity of 3.12 and 3.56. In all the top row simulations, the fibril-related parameters are set to their minimum values, with fibril radius of 80 nm, elastic modulus of 700 MPa in tension and 70 MPa in compression, and an average length of 2 µm. The bottom row shows simulation results under the same loading conditions as the top row, focusing on the impact of collagen fibril morphology and elastic modulus. In particular, to study the fibril elastic modulus effect we increase the tensile and compressive elastic moduli to 1 GPa and 100 MPa, respectively, maintaining a compression-to-tension ratio of one-tenth as in [66, 50]. These results examine inter- and intra-connectivity effects at a 4 mg/mL fixed concentration. All simulation results represent the response of ten samples per condition, with the average response shown as a line and the standard deviation indicated by the shaded area.
Figure 5 illustrates the mechanical response for various topological and microstructural parameters derived from experiments and our simulations. We examine the role of network-level factors (i.e., connectivity and concentration) and fibril-level characteristics (i.e., radius, elastic modulus, and average length) using our 3D simulations. We systematically alter these parameters within the experimentally reported ranges from [43] (Table 6.1) for diameter and average fiber length, and from [55, 30] for fibril elastic modulus. We find that transitioning from the low-stretch to the high-stretch regime results in a nonlinear stiffening for all tested topologies. In particular, for all collagen type I networks, we observe a two-order increase in stretch stiffening from the initial to the maximum differential shear modulus, with significant stiffening emerging in the medium-to-high-stretch regime. Our model results are consistent with the experimental and computational data from [43], where the concentration is fixed at 4 mg/mL while the average connectivity varies with temperature (Figure 5 - experimental data and connectivity effect). Conversely, variations in concentration exhibit an uneven impact, with the most significant changes in the differential shear modulus occurring in the low-stretch regime (Figure 5 - concentration effect). In this regime, fibrils primarily undergo non-affine rotations to align with the principal loading direction. Notably, increasing the concentration from 1.5 to 4 mg/mL results in an order-of-magnitude increase in the initial differential shear modulus. In contrast, within the large-stretch regime, the maximum modulus remains largely insensitive to concentration changes. At a concentration of 1.5 mg/mL, our models predict an almost three-order-of-magnitude increase in modulus when transitioning from the low- to high-stretch regime for both tested topologies. These results suggest that, in the nonlinear regime, network stiffness becomes independent of fiber concentration, consistent with the findings of [85, 51]. With regard to the morphological changes of the fibers, variations in radius and length made to simulate the temperature do not exhibit noticeable effects across all tested topologies and stretch ranges (Figure 5 - fibril radius and length effects). Overall, increasing the radius effect slightly increases the network’s differential shear modulus, whereas increasing the length reduces it. With the more pronounced effect in the low-stretch regime. This is because longer, more slender fibrils are more susceptible to bending, particularly in the low-stretch regime when not aligned with the primary loading direction. The limited impact arises from the narrow range of morphological variations tested, which are based on experimentally observed temperature-induced changes. These variations preserve a nearly constant bending-to-stretching stiffness ratio, therefore preventing significant changes in the network’s mechanical behavior. Increasing elastic modulus from 700 MPa to 1 GPa (Figure 5 - fibril elastic modulus effects) yields comparable behavior for the medium-to-high-stretch regime. However, it exerts a more pronounced influence in the low-stretch regime. Notably, these findings remain consistent across different connectivity conditions, suggesting that they are independent of the topology characterizing the network. Therefore, including the asymmetry in the fibril elastic modulus introduces an uneven effect in the nonlinear response of our networks. In contrast, when tension-compression asymmetry is excluded, the fiber elastic modulus directly scales with the network’s differential modulus (Figure S.3), and nonlinearity primarily arises from geometry.
@@
Figure 6 summarizes the deformation-dependent influence of each microstructural parameter on the network’s bulk elastic properties. To build the central table, we compare the variation in the differential shear modulus for each microstructural parameter by averaging data from 10 independent samples for each tested condition. At predetermined stretch points, we compute the mean and standard deviation for each condition and derive the difference between the means normalized by the combined sum of their averages and the pooled standard deviation. We aggregate these normalized changes across all parameters and compute the 40th and 70th percentiles to set thresholds: values below the 40th percentile indicate a low impact, values between the 40th and 70th percentiles indicate a moderate impact, and values above the 70th percentile indicate a high impact.
@@ 3.4 Non-affinity in simulated collagen network
Figure 7 illustrates the discrepancy between the nodal displacements predicted by our in silico simulations and the affine predictions based on the deformation gradient imposed at the boundaries.
Figure 7: Non-affinity score with nodal displacements. Comparison of predicted affine displacements (gray) and computed displacements for five samples, with average connectivity $\overline{z}=$ 3.12 (top row) and connectivity $\overline{z}=$ 3.56 (bottom row), at three distinct stretch levels. For the low-stretch regime, non-affine modes predominate, with a large set of nodes not following the boundary displacement based on the macroscopic deformation gradient. At stretch 1.5, stretch-induced affinity enhances the agreement between predicted and measured distributions.
Here, we focus on two specific cases: loosely connected networks with an average connectivity of $\overline{z}$ = 3.12, and more interconnected networks with average connectivity of $\overline{z}$ = 3.56. We define the non-affinity score ($NS$) as the normalized root mean squared error between purely affine deformation ($u_{affine}$) of the whole network (i.e., the displacements obtained by interpolating the macroscopic deformation gradient imposed at the boundaries to all internal nodes) and the displacements predicted in our simulations ($u_{predicted}$):
@@
where $n$ is the total number of nodes. The networks exhibit different behavior from the affine model under all the tested conditions. The most significant discrepancies occur in the low- to medium-stretch regime, where the $NS$ score decreases fivefold. In contrast, the medium- to high-stretch regime experiences only a twofold reduction. In the low-stretch regime, very few nodes adhere to the boundary displacement prescribed by the imposed macroscopic deformation gradient. However, a noticeable shift occurs in the medium- to high-stretch regime, where nodal displacements closely follow an overall affine displacement pattern. Both loosely connected (with $\overline{z}=$ 3.12) and more interconnected (with $\overline{z}=$ 3.56) networks converge toward a more affine response at higher stretch levels. Networks with lower connectivity exhibit slightly increased non-affinity in the low-stretch regime than those with higher connectivity. With fewer fibers aligning with the affine deformation, load distribution in these loosely connected networks becomes highly localized. Increasing the number of cross-links compared to the branches (with $\overline{z}=$ 3.56) results in a marginally improved match between the theoretical affine predictions and the experimental measurements. Overall, these networks exhibit nodal deformation patterns that, in the low- to medium-stretch regime, deviate from affine deformation assumptions.
@@ 4.1 Topological complexity and heterogeneity
Many materials - not just biological materials - are made of fibers. Natural fibrous materials include cotton, wool, and silk, while industrial examples range from polymeric foams and rubber to felt and paper. Among biological materials, an important fibrous system is the extracellular matrix, where collagen fibers assemble into networks to bear mechanical loads and transmit mechanical information over long distances [4]. Another crucial fibrous system is fibrin, derived from fibrinogen, which assembles into a fibrillar network upon tissue injury to form blood clots. These clots must withstand high tensile forces from blood flow while ensuring rapid degradation after wound healing to prevent thrombosis [77]. Accurately modeling this seemingly passive fibrous environment through computational methods remains a significant challenge. A key requirement is to account for the three-dimensional arrangement of these networks and their structural heterogeneity, which affects the local cellular environment and influences the overall mechanical response of tissues. In this study, we employ a mechanistic approach that explicitly models the interconnected fiber networks underlying soft tissues by leveraging high-resolution imaging data. This framework enables us to investigate how macroscopic material behavior emerges from its microscopic constituents, accounting for microstructural variations. Our generative pipeline builds on the framework by [53], where high-resolution confocal images are adopted to derive an image-based representation of collagen networks using simulated annealing. The iterative optimization process is designed to transform initially random Voronoi networks into structures that accurately represent the target biological network in terms of connectivity and length distribution. Compared to [53], we capture the non-affine localized effects by restricting the network to include only cross-links and branches. For the length optimization, we employ Kullback-Leibler divergence to compare the actual and target distributions throughout the iterative optimization process. As a result, we improve the energy accuracy of the optimization process without increasing runtimes compared to the commonly used [53, 52, 64, 24] Cramér-von Mises tests. Notably, while the probabilistic search strategy of simulated annealing cannot guarantee convergence of the optimization scheme to the global minimum [89], it is sufficient for our purposes that the final configuration statistically aligns with experimental measurements without requiring exact correspondence. While focused on collagen-like networks, our modeling pipeline can adapt to other biomaterials and topologies. For instance, we can effectively tune the fibers to replicate the morphology and mechanics of fibrin networks to simulate the mechanical behavior of another essential load-bearing biopolymer in the extracellular space. Alternatively, we can introduce higher connectivity by merging nodes within a specific distance threshold, or we can disregard the bending rigidity to model fibers as springs transmitting only central forces, as observed in the intracellular environment of semi-flexible F-actin networks and intermediate filaments [51]. Leveraging detailed microstructural information, our approach can be extended to model non-living networks. For example, with random Voronoi seeds, we can replicate the microstructure of paper, while with regularly arranged seeds, we can create ordered structures, such as those found in fabrics [76].
@@
4.2 Impact of microstructure on stress-stretch behavior, a case study of collagen networks